OpenStudy (anonymous):

If a chord of circle with radius 5 cm subtends a right angle at the centre , then length of chord is

OpenStudy (anonymous):

 use pythagoras theorem

OpenStudy (anonymous):

5^2 + 5^2

OpenStudy (anonymous):

\[\sqrt{25+25}=\sqrt{50}=7.071\]

OpenStudy (anonymous):

@hitten101 u forgot to put square root

OpenStudy (anonymous):

oh yes.. totally missed it